OpenStudy (anonymous):

help..... Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (anonymous):

Find Slope with points : Y2-Y1/X2-X1 Then use this : Y2-y=m(X2-x) Y1=3 X1=5 Y2=-2 X2=8 Got it ?

OpenStudy (anonymous):

yes thankx

OpenStudy (anonymous):

;)